OpenStudy (anonymous):

\[2x-6\]over\[3x-x ^{2}\]

OpenStudy (anonymous):

factor out 2 on top factor out x in the bottom

OpenStudy (anonymous):

\[2x - 6/(3x-x^2) = 2(x-3)/(-x)(x-3) = -2/x\]